OpenStudy (anonymous):

Length Width Rectangle A 2x + 8 x - 1 Rectangle B 3x + 5 2x - 2 WHAT IS THE RESULT OF (PERIMETER OF B) MINUS (PERIMETER OF A)

OpenStudy (anonymous):

Hey BasedGod1122 again. Do you remember how to find the perimeter of a rectangle?

OpenStudy (anonymous):

Yes of course. The length (x2) plus the width (x2) Which is just adding up the sides.

OpenStudy (anonymous):

That's correct. So first let's find the perimeter of rectangle A. P = 2L + 2W P = 2(2x+8) + 2(x-1) Do you get it?

OpenStudy (anonymous):

(4x+16) + (2x-2) correct?

OpenStudy (anonymous):

yes correct!

OpenStudy (anonymous):

and then just combine like terms

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

6x+14?

OpenStudy (anonymous):

Yeah that's the perimeter of rectangle A.

OpenStudy (anonymous):

So now try to do the same thing for the perimeter of rectangle B.

OpenStudy (anonymous):

P = 2(3x+5) + 2(2x-2)

OpenStudy (anonymous):

(6x + 10) (4x-4)

OpenStudy (anonymous):

Yes correct but you forgot to add the addition sign. It would look like this: (6x+10)+(4x-4)

OpenStudy (anonymous):

Combine like terms again

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

10x and 6?

OpenStudy (anonymous):

10x + 6 and that's the perimeter of rectangle B

OpenStudy (anonymous):

So what the question is asking (perimeter of B) - (perimeter of A) so just plug in the what you got. (10x +6) - (6x + 14)

OpenStudy (anonymous):

Remember to distribute the negative sign first.

OpenStudy (anonymous):

And then combine like terms

OpenStudy (anonymous):

4x -8

OpenStudy (anonymous):

Yes that's correct. Is that one of the answers?

OpenStudy (anonymous):

Yes. thank you again turquoisee

OpenStudy (anonymous):

You're welcome! :)